OpenStudy (anonymous):

57 is what percent of 60?

OpenStudy (gorv):

let 57 is x percent of 60 57=60*x/100

OpenStudy (gorv):

calculate the value pof x from here

OpenStudy (anonymous):

Before I looked at your answer I put 57/60 and got 95

OpenStudy (gorv):

\[57=60*\frac{ x }{ 100 }\]

OpenStudy (texaschic101):

there is a formula for this...the percent formula.. is/of = %/100 57 is what percent of 60.. "is" = 57 % = x "of " = 60 now sub into the formula 57/60 = x/100 -- this is a proportion, so cross multiply (60)(x) = (57)(100) 60x = 5700 x = 5700/60 x = 95 check.. 57 is 95% of 60 57 = 0.95(60) 57 = 57 (correct) your answer is 95%

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

No, thank you both

OpenStudy (texaschic101):

your quite welcome :)

OpenStudy (paki):

57 is what percent of 60? 57 = x /100 * 60 or use this... is equals to = what equals to x percent equals to /100 of equals to multiply sign

OpenStudy (texaschic101):

that works just as well....I guess it is what you are most comfortable doing..either way will work

OpenStudy (paki):

agree with @texaschic101

OpenStudy (texaschic101):

:)